# multiple directions/.style=for tree=#1, phantom, for relative level=1no edge, delay=!c.content/.pgfmath=content("!u"), before computing xy=l=0,s=0, multiple directions/.default=, grow subtree/.style=for tree=grow=#1, grow’ subtree/.style=for tree=grow’=#1

### v1

@@ 4.2.3 Proof of the technical lemmas
We start by proving a stronger version of Lemma 19, where we show the asymptotic correspondence not just for $w_{i,j}$ but for a well-behaved local function.
@@
Let $f:\hat{\mathcal{G}}^{*}\mapsto\mathbb{R}$. We say that $f$ is a local function if there exists some $H>0$ such that for any $T,T^{\prime}\in\hat{\mathcal{G}}^{*}$,
$N_{H}(T)=N_{H}(T^{\prime})\Rightarrow f(T)=f(T^{\prime}).$
Proposition 23.
@@
Lemma 19 follows by taking $f(G,o)=w(o)$ that is local with $H=0$.
Proof of Proposition 23.
@@
$\mathbb{E}\left[\frac{1}{|\overset{\rightarrow}{E}_{n}|}\sum_{(i,j)\in V_{n}^{2}}q_{i,j,H,n,x}f_{i,j}\right]=\int_{\hat{\mathcal{G}}^{*}}g_{H}(T)\mathbbm{1}_{w(o)\leq x}f(T)\mathrm{d}\hat{\lambda}_{n,H}(T)$
Let $M$ be the Lipschitz constant of $f$ on the weights of $N_{H}(T)$. As $f\mathbbm{1}_{w(o)\leq x}$ is bounded by $Mx$ and $g_{H}$ by $1$, we have that
$\int_{\hat{\mathcal{G}}^{*}}g_{H}(T)\mathbbm{1}_{w(o)\leq x}f(T)\mathrm{d}\hat{\lambda}_{n,H}(T)\underset{n\rightarrow\infty}{\longrightarrow}\int_{\hat{\mathcal{G}}^{*}}g_{H}(T)\mathbbm{1}_{w(o)\leq x}f(T)\mathrm{d}\hat{\lambda}_{H}(T).$
@@
$\mathbb{E}\left[\frac{2|E_{n}|}{|V_{n}|}\frac{1}{|\overset{\rightarrow}{E}_{n}|}\sum_{(i,j)\in V_{n}^{2}}q_{i,j,H,n,x}f_{i,j}\mathbbm{1}_{(i,j)\in E_{n}}\right]\underset{n\rightarrow\infty}{\longrightarrow}\mathbb{E}_{N\sim\pi}[N]\int_{\hat{\mathcal{G}}^{*}}g_{H}(T)\mathbbm{1}_{w(o)\leq x}f(T)\mathrm{d}\hat{\lambda}_{H}(T)$
@@
$\int_{\hat{\mathcal{G}}^{*}}g_{H}(T)\mathbbm{1}_{w(o)\leq x}f(T)\mathrm{d}\hat{\lambda}_{H}(T)\underset{n\rightarrow\infty}{\longrightarrow}\int_{\hat{\mathcal{G}}^{*}}g_{H}(T)f(T)\mathrm{d}\hat{\lambda}_{H}(T).$
@@
$\int_{\hat{\mathcal{G}}^{*}}g_{H}(T)f(T)\mathrm{d}\hat{\lambda}_{H}(T)=\int_{\hat{\mathcal{G}}^{*}}g(T)f(T)\mathrm{d}\hat{\lambda}(T)=\mathbb{E}[f(\mathbb{T},o)\mathbbm{1}_{o\in\mathbb{M}}].$
@@
by Proposition 6. ∎
Proof of Lemma 20.
@@
$\mathbb{E}\left[\frac{1}{|V_{n}|}\sum_{j\in V_{n}}\left|q_{j,j,H,n,x}\right|_{-}\right]=\int_{\mathcal{G}^{*}}\left|1-\sum_{v\sim o}g_{H}(T,(o,v))\mathbbm{1}_{w(o)\leq x}\right|_{-}\mathrm{d}\lambda_{n,H}(T).$
@@
$\int_{\mathcal{G}^{*}}\left|1-\sum_{v\sim o}g_{H}(T,(o,v))\mathbbm{1}_{w(o)\leq x}\right|_{-}\mathrm{d}\lambda_{n,H}(T)\leq\int_{\mathcal{G}^{*}}\left|1-\sum_{v\sim o}g_{H}(T,(o,v))\right|_{-}\mathrm{d}\lambda_{n,H}(T).$
Now with a proof almost identical to the one of Lemma 18, we can prove that for any $H$, $\lambda_{n,H}$ converges weakly to $\lambda_{H}$. Since $\mathbb{E}_{N\sim\pi}[N]<\infty$, we can first integrate on $\deg(o)$ to bound the integrand by $\mathbb{E}_{N\sim\pi}[N]$. So as $n\rightarrow\infty$, we get:
$\int_{\mathcal{G}^{*}}\left|1-\sum_{v\sim o}g_{H}(T,(o,v))\right|_{-}\mathrm{d}\lambda_{H}(T)$
@@
$\sum_{v\sim o}g(T,(o,v))=\mathbb{P}\left(o\text{ is matched by }\mathbb{M}|\mathbb{T}=T\right)\leq 1\,\,\,\,\lambda-\text{almost surely on }\mathcal{G}^{*}$
@@
$\int_{\mathcal{G}^{*}}\left|1-\sum_{v\sim o}g_{H}(T,(o,v))\right|_{-}\mathrm{d}\lambda(T)=0.$
Proof of Proposition 21.
@@
Thus, we set $m_{k_{i},j_{i}^{\prime}}^{(t+1)}\leftarrow m_{k_{i},j_{i}^{\prime}}^{(t+1)}-\alpha^{(t+1)}$ and $L_{k_{i}}^{(t+1)}\leftarrow L_{k_{i}}^{(t+1)}-\alpha^{(t+1)}$ for $1\leq i\leq p$ , $m_{k_{i},j_{i}^{\prime}}^{(+1)}\leftarrow m_{k_{i},j_{i}^{\prime}}^{(+1)}+\frac{p}{l+1}\alpha^{(t+1)}$ and $L_{k_{i}}^{(t+1)}\leftarrow L_{k_{i}}^{(t+1)}+\frac{p}{l+1}\alpha^{(t+1)}$ for $n-l\leq i\leq n$.
@@
We then apply the same algorithm to the columns in order to obtain a matrix $\tilde{M}$ whose rows and columns sum to the same amount $\hat{L}$, which is $\hat{L}:=\frac{1}{n}\sum_{i,j}m_{i,j}$. Finally we output $S:=\frac{1}{\hat{L}}\tilde{M}$ that is a bistochastic matrix.
@@
$|\hat{L}-1|$ $=\left|\frac{1}{n}\sum_{i}L_{i}-1\right|=\frac{1}{n}\left|\sum_{i}L_{i}-\sum_{i}\frac{1}{n}\right|$
$\leq\frac{1}{n}\sum_{i}\left|L_{i}-1\right|$
@@
$<\sum_{i}|L_{i}-1|+\sum_{j}|C_{j}-1|+2n|1-\hat{L}|$
$<n\varepsilon+2n\varepsilon=3n\varepsilon.$
For (29), since $\frac{1}{2}>\varepsilon>0$, we have $|1-\frac{1}{\hat{L}}|=|\frac{1-\hat{L}}{\hat{L}}|<4\varepsilon$. Finally, we have that $n\hat{L}=\sum_{i,j}\tilde{m}_{i,j}=\sum_{i,j}m_{i,j}^{(0)}<n(1+\varepsilon)$.
@@
$\|S-M\|_{1}<3\varepsilon+4n\varepsilon(1+\varepsilon)<12n\varepsilon$
@@
$\leq d(M,\text{Sym}_{n})+\inf_{S\text{ bistochastic symmetric}}\left\|\frac{M+M^{T}}{2}-\frac{S+S^{T}}{2}\right\|_{1}.$
Now take $S^{*}$ such that $d(M,\text{Birk}_{n})=\|M-S^{*}\|_{1}$, then $\frac{S^{*}+S^{*T}}{2}$ is bistochastic symmetric so:
$\inf_{S\text{ bistochastic symmetric}}\|M-S\|_{1}$ $\leq d(M,\text{Sym}_{n})+\inf_{S\text{ bistochastic symmetric}}\left\|\frac{M+M^{T}}{2}-\frac{S+S^{T}}{2}\right\|_{1}$
$\leq d(M,\text{Sym}_{n})+\left\|\frac{M+M^{T}}{2}-\frac{S^{*}+S^{*T}}{2}\right\|_{1}$
$\leq d(M,\text{Sym}_{n})+\left\|\frac{M-S^{*}}{2}\right\|_{1}+\left\|\frac{M^{T}-S^{*T}}{2}\right\|_{1}$
$\leq d(M,\text{Sym}_{n})+d(M,\text{Birk}_{n}).$
5 Proof of the main results
This Section is devoted to the remaining of the proof of Theorem 1 and to the proof of Theorem 2.
5.1 Convergence of optimally matched graphs (Theorem 1)
Let us conclude the proof of Theorem 1 by combining the main results of Section 3 and Section 4.
Section 4 shows that every subsequence of $(G_{n},o_{n},M_{\mathrm{opt}}(G_{n}))$ has a subsequence that converges locally to some $(\mathbb{T},\mathbb{M})$ where $\mathbb{M}$ is an optimal matching on $\mathbb{T}$. Section 3 shows that the only optimal matching in law on the UBGW tree $\mathbb{T}$ is $\mathbb{M}_{h}$. So every subsequence of $(G_{n},o,M_{\mathrm{opt}}(G_{n}))$ has a subsequence that converges to $(\mathbb{T},\mathbb{M}_{h})$. We deduce that $(G_{n},o,M_{\mathrm{opt}}(G_{n}))$ converges locally to $(\mathbb{T},\mathbb{M}_{h})$. This concludes the proof of Theorem 1.
5.2 Uniqueness of the message passing distribution (Theorem 2)
We now turn to the proof of Theorem 2. The idea is to recover the law of $\zeta_{h}$ from families of statistics on $\mathbb{M}_{h}$.
@@ 5.2 Uniqueness of the message passing distribution (Theorem 2)
$\mathbb{P}\left(w(o)\in\mathbb{M}|w(o)=x\right)$
@@
Evaluated for $\mathbb{M}_{h_{1}}$ and $\mathbb{M}_{h_{2}}$, this yields:
$\mathbb{P}\left(Z_{h_{1}}+Z_{h_{1}}^{\prime}<x\right)=\mathbb{P}\left(Z_{h_{2}}+Z_{h_{2}}^{\prime}<x\right),$
where $Z_{h_{1}},Z_{h_{2}},Z_{h_{1}}^{\prime},Z_{h_{2}}^{\prime}$ are independent variables, and the cdf of $Z_{h_{1}},Z_{h_{1}}^{\prime}$ is $h_{1}$, the cdf of $Z_{h_{2}},Z_{h_{2}}^{\prime}$ is $h_{2}$. This shows that for any $t\in\mathbb{R}$:
$\mathbb{E}\left[e^{itZ_{h_{1}}}\right]^{2}=\mathbb{E}\left[e^{it(Z_{h_{1}}+Z_{h_{1}}^{\prime})}\right]=\mathbb{E}\left[e^{it(Z_{h_{2}}+Z_{h_{2}}^{\prime})}\right]=\mathbb{E}\left[e^{itZ_{h_{2}}}\right]^{2}$
[…]
@@
Let us assume that $\hat{\pi}_{1}>0$, we will condition on the event that the weight of the root edge is $w_{0}$ and that the $+$ side of the root is a simple path $v_{1},v_{2},....,v_{H}$ of length $H$ of weights $w_{1},....,w_{H}$. We will further condition on the event that $v_{H}$ is unmatched in the matching $\mathbb{M}_{h}$. The total event of conditioning on the sequence of weights and $v_{H}$ being unmatched is measurable in the matching and graph, Figure 16 gives a depiction of the situation.
@@
We will show in Corollary 3 that $\mathbb{P}(Z^{\prime}=0)=h(0)$ is invariant for $h=h_{1}$ or $h=h_{2}$ , which then implies that:
@@
is invariant for $h=h_{1}$ or $h=h_{2}$. The probability in the previous display can be written as the following integral:
$G(w_{i}):=\int_{\mathbb{R}_{+}}\mathbbm{1}_{\left(w_{H}\leq\max\left(0,w_{H-1}-\max\left(0,w_{H-2}-....-\max\left(0,w_{1}-\max\left(0,w_{0}-x\right)\right)...\right)\right)\right)}\mathrm{d}\mathbb{P}_{h}(x).$
@@
$g(x,w_{i}):=g(x_{,}w_{0},...,w_{H})=\mathbbm{1}_{\left(w_{H}\leq\max\left(0,w_{H-1}-\max\left(0,w_{H-2}-....-\max\left(0,w_{1}-\max\left(0,w_{0}-x\right)\right)...\right)\right)\right)},$
@@
$G(w_{i})=\int_{\mathbb{R}_{+}}g(x,w_{i})\mathrm{d}\mathbb{P}_{h}(x).$
To prove Theorem 2, we will show that for any $x_{0}\in[0,\sup(\mathrm{supp}(\omega))]$, for $\varepsilon>0$ small enough, we can recover:
$\mathbb{P}(x_{0}\leq Z<x_{0}+\varepsilon)$
as a difference $G(w_{i})-G(w_{i}^{\prime})$ for $(w_{i})$ and $(w_{i}^{\prime})$ chosen according to the following Lemma, which suffices to recover the law of $h$.
@@
Figure 17: Illustration of Lemma 24 and its consequences.
Lemma 24.
@@
    The sequence $\overline{u}_{k}:=\sum_{i=0}^{2k}(-1)^{i}x_{i}$ is positive decreasing.
@@
    The sequence $\underline{u}_{k}:=\sum_{i=0}^{2k+1}(-1)^{i}x_{i}$ is increasing.
@@
    Finally, $\overline{u}_{H}>\underline{u}_{H}$.
Proof of Lemma 24.
@@
We are now ready to prove the uniqueness in law of the variables $Z$:
Proof of Theorem 2.
By assumption on $\omega$ take $b=\sup(\mathrm{supp}(\omega))$ and $a$ such that $[a,b]\subset\mathrm{supp}(\omega)$. Now for any $h$, take any $x_{0}\in[0,b]$ and take $H>0$ and a sequence $(w_{0},....,w_{H})$ given by Lemma 24. We will reuse the notations $\overline{u}_{k}$ and $\underline{u}_{k}$ from Lemma 24. We will only treat the case when $x_{0}>\frac{b}{2}$ and $H$ even, the other case being similar.
Recall $g(x,w_{0},...,w_{H})=g(x_{,}w_{i})=\mathbbm{1}_{\left(w_{H}\leq\max\left(0,w_{H-1}-\max\left(0,w_{H-2}-....-\max\left(0,w_{1}-\max\left(0,w_{0}-x\right)\right)...\right)\right)\right)}$. Decompose the invariant integral we found earlier:
$\int_{\mathbb{R}_{+}}g(x,w_{i})\mathrm{d}\mathbb{P}_{h}(x)$ $=\int_{0}^{\underline{u}_{0}}g(x,w_{i})\mathrm{d}\mathbb{P}_{h}(x)+\sum_{k=0}^{\frac{H}{2}-1}\int_{\underline{u}_{k}}^{\underline{u}_{k+1}}g(x,w_{i})\mathrm{d}\mathbb{P}_{h}(x)$
$\qquad+\int_{\underline{u}_{\frac{H}{2}-1}}^{\overline{u}_{\frac{H}{2}}}g(x,w_{i})\mathrm{d}\mathbb{P}_{h}(x)$
$\qquad+\sum_{k=0}^{\frac{H}{2}-1}\int_{\overline{u}_{\frac{H}{2}-k}}^{\overline{u}_{\frac{H}{2}-1-k}}g(x,w_{i})\mathrm{d}\mathbb{P}_{h}(x)+\int_{\overline{u}_{0}}^{b}g(x,w_{i})\mathrm{d}\mathbb{P}_{h}(x).$
@@
    $\forall x\in\left(\underline{u}_{\frac{H}{2}-1},\overline{u}_{\frac{H}{2}}\right),\,g(x,w_{i})=0.$
@@
    $\forall x\in\left[\overline{u}_{\frac{H}{2}},\overline{u}_{\frac{H}{2}-1}\right),g(x,w_{i})=1.$
@@
    then none of the $g(x,w_{i})$’s value would change on the other intervals.
@@
Looking at the inequality for $\overline{u}_{0}=w_{0}>x$ it means that the last nested $\max(0,w_{0}-x)$ intervening in $g(x,w_{i})$ evaluates at $w_{0}-x$.

### v2

@@ 4.2.3 Proof of the technical lemmas
We start by proving a stronger version of Lemma 4.5, where we show the asymptotic correspondence not just for $w_{i,j}$ but for a well-behaved local function.
@@
Let $f:\hat{\mathcal{G}}^{\star}\mapsto\mathbb{R}$. We say that $f$ is a local function if there exists some $H>0$ such that for any $T,T^{\prime}\in\hat{\mathcal{G}}^{\star}$,
$B_{H}(T)=B_{H}(T^{\prime})\Rightarrow f(T)=f(T^{\prime}).$
Proposition 4.10.
@@
Lemma 4.5 follows by taking $f(G,o)=w(o)$ that is local with $H=0$.
Proof of Proposition 4.10.
@@
$\mathbb{E}\left[\frac{1}{|\overset{\rightarrow}{E}_{n}|}\sum_{(i,j)\in V_{n}^{2}}q_{i,j,H,n,x}f_{i,j}\right]=\int_{\hat{\mathcal{G}}^{\star}}g_{H}(T)\mathbbm{1}_{w(o)\leq x}f(T)\mathrm{d}\hat{\lambda}_{n,H}(T)$
Let $M$ be the Lipschitz constant of $f$ on the weights of $B_{H}(T)$. As $f\mathbbm{1}_{w(o)\leq x}$ is bounded by $Mx$ and $g_{H}$ by $1$, we have that
$\int_{\hat{\mathcal{G}}^{\star}}g_{H}(T)\mathbbm{1}_{w(o)\leq x}f(T)\mathrm{d}\hat{\lambda}_{n,H}(T)\underset{n\rightarrow\infty}{\longrightarrow}\int_{\hat{\mathcal{G}}^{\star}}g_{H}(T)\mathbbm{1}_{w(o)\leq x}f(T)\mathrm{d}\hat{\lambda}_{H}(T).$
@@
$\mathbb{E}\left[\frac{2|E_{n}|}{|V_{n}|}\frac{1}{|\overset{\rightarrow}{E}_{n}|}\sum_{(i,j)\in V_{n}^{2}}q_{i,j,H,n,x}f_{i,j}\mathbbm{1}_{(i,j)\in E_{n}}\right]\underset{n\rightarrow\infty}{\longrightarrow}\mathbb{E}_{N\sim\pi}[N]\int_{\hat{\mathcal{G}}^{\star}}g_{H}(T)\mathbbm{1}_{w(o)\leq x}f(T)\mathrm{d}\hat{\lambda}_{H}(T)$
@@
$\int_{\hat{\mathcal{G}}^{\star}}g_{H}(T)\mathbbm{1}_{w(o)\leq x}f(T)\mathrm{d}\hat{\lambda}_{H}(T)\underset{n\rightarrow\infty}{\longrightarrow}\int_{\hat{\mathcal{G}}^{\star}}g_{H}(T)f(T)\mathrm{d}\hat{\lambda}_{H}(T).$
@@
$\int_{\hat{\mathcal{G}}^{\star}}g_{H}(T)f(T)\mathrm{d}\hat{\lambda}_{H}(T)=\int_{\hat{\mathcal{G}}^{\star}}g(T)f(T)\mathrm{d}\hat{\lambda}(T)=\mathbb{E}[f(\mathbb{T},o)\mathbbm{1}_{o\in\mathbb{M}}].$
@@
by Proposition 2.12. ∎
Proof of Lemma 4.6.
@@
$\mathbb{E}\left[\frac{1}{|V_{n}|}\sum_{j\in V_{n}}\left|q_{j,j,H,n,x}\right|_{-}\right]=\int_{\mathcal{G}^{\star}}\left|1-\sum_{v\sim o}g_{H}(T,(o,v))\mathbbm{1}_{w(o)\leq x}\right|_{-}\mathrm{d}\lambda_{n,H}(T).$
@@
$\int_{\mathcal{G}^{\star}}\left|1-\sum_{v\sim o}g_{H}(T,(o,v))\mathbbm{1}_{w(o)\leq x}\right|_{-}\mathrm{d}\lambda_{n,H}(T)\leq\int_{\mathcal{G}^{\star}}\left|1-\sum_{v\sim o}g_{H}(T,(o,v))\right|_{-}\mathrm{d}\lambda_{n,H}(T).$
Now with a proof almost identical to the one of Lemma 4.3, we can prove that for any $H$, $\lambda_{n,H}$ converges weakly to $\lambda_{H}$. Since $\mathbb{E}_{N\sim\pi}[N]<\infty$, we can first integrate on $\deg(o)$ to bound the integrand by $\mathbb{E}_{N\sim\pi}[N]$. So as $n\rightarrow\infty$, we get:
$\int_{\mathcal{G}^{\star}}\left|1-\sum_{v\sim o}g_{H}(T,(o,v))\right|_{-}\mathrm{d}\lambda_{H}(T)$
@@
$\sum_{v\sim o}g(T,(o,v))=\mathbb{P}\left(o\text{ is matched by }\mathbb{M}|\mathbb{T}=T\right)\leq 1\,\,\,\,\lambda-\text{almost surely on }\mathcal{G}^{\star}$
@@
$\int_{\mathcal{G}^{\star}}\left|1-\sum_{v\sim o}g_{H}(T,(o,v))\right|_{-}\mathrm{d}\lambda(T)=0.$
Proof of Proposition 4.7.
@@
Thus, we set $m_{k_{i},j_{i}^{\prime}}^{(t+1)}\leftarrow m_{k_{i},j_{i}^{\prime}}^{(t+1)}-\alpha^{(t+1)}$ and $L_{k_{i}}^{(t+1)}\leftarrow L_{k_{i}}^{(t+1)}-\alpha^{(t+1)}$ for $1\leq i\leq p$ , $m_{k_{i},j_{i}^{\prime}}^{(+1)}\leftarrow m_{k_{i},j_{i}^{\prime}}^{(+1)}+\frac{p}{l+1}\alpha^{(t+1)}$ and $L_{k_{i}}^{(t+1)}\leftarrow L_{k_{i}}^{(t+1)}+\frac{p}{l+1}\alpha^{(t+1)}$ for $d-l\leq i\leq d$.
@@
We then apply the same algorithm to the columns in order to obtain a matrix $\tilde{M}$ whose rows and columns sum to the same amount $\hat{L}$, which is $\hat{L}:=\frac{1}{d}\sum_{i,j}m_{i,j}$. Finally we output $S:=\frac{1}{\hat{L}}\tilde{M}$ that is a bistochastic matrix.
@@
$|\hat{L}-1|$ $=\left|\frac{1}{d}\sum_{i}L_{i}-1\right|=\frac{1}{d}\left|\sum_{i}L_{i}-\sum_{i}\frac{1}{d}\right|$
$\leq\frac{1}{d}\sum_{i}\left|L_{i}-1\right|$
@@
$<\sum_{i}|L_{i}-1|+\sum_{j}|C_{j}-1|+2d|1-\hat{L}|$
$<d\varepsilon+2d\varepsilon=3d\varepsilon.$
For (35), since $\frac{1}{2}>\varepsilon>0$, we have $|1-\frac{1}{\hat{L}}|=|\frac{1-\hat{L}}{\hat{L}}|<4\varepsilon$. Finally, we have that $d\hat{L}=\sum_{i,j}\tilde{m}_{i,j}=\sum_{i,j}m_{i,j}^{(0)}<d(1+\varepsilon)$.
@@
$\|S-M\|_{1}<3d\varepsilon+4d\varepsilon(1+\varepsilon)<12d\varepsilon$
@@
$\leq d(M,\text{Sym}_{d})+\inf_{S\text{ bistochastic symmetric}}\left\|\frac{M+M^{T}}{2}-\frac{S+S^{T}}{2}\right\|_{1}.$
Now take $S^{\star}$ such that $d(M,\text{Birk}_{n})=\|M-S^{\star}\|_{1}$, then $\frac{S^{\star}+S^{\star T}}{2}$ is bistochastic symmetric so:
$\inf_{S\text{ bistochastic symmetric}}\|M-S\|_{1}$ $\leq d(M,\text{Sym}_{d})+\inf_{S\text{ bistochastic symmetric}}\left\|\frac{M+M^{T}}{2}-\frac{S+S^{T}}{2}\right\|_{1}$
$\leq d(M,\text{Sym}_{d})+\left\|\frac{M+M^{T}}{2}-\frac{S^{\star}+S^{\star T}}{2}\right\|_{1}$
$\leq d(M,\text{Sym}_{d})+\left\|\frac{M-S^{\star}}{2}\right\|_{1}+\left\|\frac{M^{T}-S^{\star T}}{2}\right\|_{1}$
$\leq d(M,\text{Sym}_{d})+d(M,\text{Birk}_{d}).$
4.3 Combining everything: convergence of optimally matched graphs
Proposition 4.1 shows that every subsequence of $(G_{n},o_{n},M_{\mathrm{opt}}(G_{n}))$ has a subsequence that converges locally to some $(\mathbb{T},\mathbb{M})$ where $\mathbb{M}$ is an optimal matching on $\mathbb{T}$. Meanwhile, Theorem 1 states that the only optimal matching in law on the UBGW tree $\mathbb{T}$ is $\mathbb{M}_{h}$. So every subsequence of $(G_{n},o_{n},M_{\mathrm{opt}}(G_{n}))$ has a subsequence that converges to $(\mathbb{T},\mathbb{M}_{h})$. We deduce that $(G_{n},o_{n},M_{\mathrm{opt}}(G_{n}))$ converges locally to $(\mathbb{T},\mathbb{M}_{h})$. This concludes the proof of Theorem 2.
5 Uniqueness of the message passing distribution : Theorem 3
@@
As a warm-up, we first inspect the simple case where $\mathrm{supp}(\omega)=\mathbb{R}_{+}$. We will look at the probability that the root is matched conditioned on the weight being $x\in\mathbb{R}_{+}$:
$\mathbb{P}\left(o\in\mathbb{M}|w(o)=x\right)$
@@
From the edge-decision rule of Proposition 3.2,
$\mathbb{P}\left(Z_{h}+Z_{h}^{\prime}<x\right)=\mathbb{P}\left(o\in\mathbb{M}|w(o)=x\right)$
where $Z_{h},Z_{h}^{\prime}$ are independent variables of distribution $\zeta_{h}$. Hence, this quantity does not depend on $h$. This shows that for any $t\in\mathbb{R}$:
$\mathbb{E}\left[e^{itZ_{h}}\right]^{2}$
does not depend on $h$. By continuity the characteristic function is uniquely determined in a neighbourhood of 00, and outside of pathological cases, must determine the law of $Z_{h}$.
@@
Proof.
Let us assume that $\hat{\pi}_{1}>0$, we will condition on the event that the weight of the root edge is $w_{0}$ and that the $+$ side of the root is a simple path $v_{1},v_{2},....,v_{H}$ of length $H$ of weights $w_{1},....,w_{H}$. We will compute the probability that $v_{H}$ is unmatched in the matching $\mathbb{M}_{h}$. The total event of conditioning on the sequence of weights and $v_{H}$ being unmatched is measurable in the matching and graph, Figure 16 gives a depiction of the situation.
@@
By independence we can write that the probability of this event is the product
$\mathbb{P}(Z^{\prime}=0)\mathbb{P}(w_{h}<Z_{h}(v_{H},v_{H-1}))$
which does not depend on $h$. We will show in Corollary 1 that $\mathbb{P}(Z^{\prime}=0)=h(0)$ only depends on $\mathbb{M}_{h}$ , which then implies that:
@@
does not depend on $h$. The probability in the previous display can be written as the following integral:
$G(w_{0:H}):=G(w_{0},...,w_{H})=\int_{\mathbb{R}_{+}}\mathbbm{1}_{\left(w_{H}\leq\max\left(0,w_{H-1}-\max\left(0,w_{H-2}-....-\max\left(0,w_{1}-\max\left(0,w_{0}-x\right)\right)...\right)\right)\right)}\mathrm{d}\mathbb{P}_{h}(x).$
@@
$g(x,w_{0:H}):=g(x_{,}w_{0},...,w_{H})=\mathbbm{1}_{\left(w_{H}\leq\max\left(0,w_{H-1}-\max\left(0,w_{H-2}-....-\max\left(0,w_{1}-\max\left(0,w_{0}-x\right)\right)...\right)\right)\right)},$
@@
$G(w_{0:H})=\int_{\mathbb{R}_{+}}g(x,w_{0:H})\mathrm{d}\mathbb{P}_{h}(x).$
To prove Theorem 3, we will show that for any $x_{0}\in[0,\sup(\mathrm{supp}(\omega))]$, for $\varepsilon>0$ small enough, we can recover $\mathbb{P}(x_{0}\leq Z<x_{0}+\varepsilon)$ as a difference $G(w_{0:H})-G(w_{0:H}^{\prime})$ for $(w_{0:H})$ and $(w_{0:H}^{\prime})$ with $(w_{0:H})$ chosen according to the following Lemma and $(w_{0:H}^{\prime})=(w_{0},...,w_{H-1},w_{H}+\varepsilon)$. ∎
@@
Figure 17: Illustration of Lemma 5.1 and its consequences.
Lemma 5.1.
@@
    The sequence $\overline{u}_{k}:=\sum_{i=0}^{2k}(-1)^{i}w_{i}$ is positive decreasing.
@@
    The sequence $\underline{u}_{k}:=\sum_{i=0}^{2k+1}(-1)^{i}w_{i}$ is increasing.
@@
    Finally, $\overline{u}_{\lfloor\frac{H}{2}\rfloor}>\underline{u}_{\lfloor\frac{H}{2}\rfloor}$.
Proof of Lemma 5.1.
@@
We return to the proof of Theorem 3.
Continuation of proof of Theorem 3.
By assumption on $\omega$ take $b=\sup(\mathrm{supp}(\omega))$ and $a$ such that $[a,b]\subset\mathrm{supp}(\omega)$. Now for any $h$, take any $x_{0}\in[0,b[$ and take $H>0$ and a sequence $(w_{0},....,w_{H})$ given by Lemma 5.1. If $b=+\infty$ then do the same reasoning with an arbitrary large $b>0$. We will reuse the notations $\overline{u}_{k}$ and $\underline{u}_{k}$ from Lemma 5.1. We will only treat the case when $x_{0}>\frac{b}{2}$ and $H$ even, the other case being similar.
Recall $g(x,w_{0},...,w_{H})=g(x_{,}w_{0:H})=\mathbbm{1}_{\left(w_{H}\leq\max\left(0,w_{H-1}-\max\left(0,w_{H-2}-....-\max\left(0,w_{1}-\max\left(0,w_{0}-x\right)\right)...\right)\right)\right)}$. Decompose the invariant integral we found earlier:
$\int_{\mathbb{R}_{+}}g(x,w_{0:H})\mathrm{d}\mathbb{P}_{h}(x)$ $=\int_{0}^{\underline{u}_{0}}g(x,w_{0:H})\mathrm{d}\mathbb{P}_{h}(x)+\sum_{k=0}^{\frac{H}{2}-1}\int_{\underline{u}_{k}}^{\underline{u}_{k+1}}g(x,w_{0:H})\mathrm{d}\mathbb{P}_{h}(x)$
$+\int_{\underline{u}_{\frac{H}{2}-1}}^{\overline{u}_{\frac{H}{2}}}g(x,w_{0:H})\mathrm{d}\mathbb{P}_{h}(x)$
$+\sum_{k=0}^{\frac{H}{2}-1}\int_{\overline{u}_{\frac{H}{2}-k}}^{\overline{u}_{\frac{H}{2}-1-k}}g(x,w_{0:H})\mathrm{d}\mathbb{P}_{h}(x)+\int_{\overline{u}_{0}}^{b}g(x,w_{0:H})\mathrm{d}\mathbb{P}_{h}(x).$
@@
    $\forall x\in\left(\underline{u}_{\frac{H}{2}-1},\overline{u}_{\frac{H}{2}}\right),\,g(x,w_{0:H})=0.$
@@
    $\forall x\in\left[\overline{u}_{\frac{H}{2}},\overline{u}_{\frac{H}{2}-1}\right),g(x,w_{0:H})=1.$
@@
    then none of the $g(x,w_{0:H})$’s value would change on the other intervals.
@@
Looking at the inequality for $\overline{u}_{0}=w_{0}>x$ it means that the last nested $\max(0,w_{0}-x)$ intervening in $g(x,w_{0:H})$ evaluates at $w_{0}-x$.